OpenStudy (anonymous):

Solve and check each of the equations: x^2-4x+3=0 x^2+6x+5=0 4-x(x-3)=0 3x^2-5x=36-2x 2x(x+1) =12

OpenStudy (anonymous):

\[x1 and x2=\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\] when \[ax ^{2}+bx+c=0\]

OpenStudy (anonymous):

but you can often use a quick simpler method. find the factors of c that add to give b so for example the first equation \[x^2 -4x +3 = 0\] factors of +3 are -1 and -3 they add to give -4 so we know to use those two factors like such \[(x-1) (x-3) \] and thats the first question factored the easy way.

OpenStudy (anonymous):

so for the equation to equal zero x = 1 or x = 3, check it by using the 1 and 3 in the original equation and if it's zero , then all good :)

OpenStudy (anonymous):

try the second one now and see how you go ?